# Mamba Can Learn Low-Dimensional Targets In-Context via Test-Time Feature Learning

### v2

@@ 1.2 Related Works
Low-dimensional target function classes, such as sparse parities (Barak et al., 2022; Suzuki et al., 2023; Glasgow, 2024), signal-noise models (Allen-Zhu and Li, 2020; Cao et al., 2022), are widely adopted as theoretical benchmarks for studying a neural network’s ability to perform feature learning. This work specifically focuses on the single-index model. A line of theoretical work has analyzed the learning of these models and has established key results on sample complexity. The required sample complexity is governed by either the information exponent (for algorithms utilizing correlational information (Arous et al., 2021; Bietti et al., 2022; Damian et al., 2023; Mousavi-Hosseini et al., 2023)) or the generative exponent (for algorithms that employ suitable label transformations (Damian et al., 2024; Lee et al., 2024; Arnaboldi et al., 2024; Joshi et al., 2024)). We discuss these sample complexity results in more detail in Section 3.1.
@@ 2.1 Data Distribution for In-Context Learning
    We draw the feature vector ${\bm{\beta}}\in\mathbb{R}^{d}$ from the intrinsic feature space $S_{r}$, which is the unit sphere of a low-dimensional subspace with dimension $r$. This subspace is defined as:
    ${\bm{\beta}}\sim\mathrm{Unif}(S_{r}),\quad\text{where }S_{r}:=\left\{\bm{\theta}\in\mathbb{R}^{d}:\left\|\bm{\theta}\right\|=1,\bm{\theta}[j]=0\text{ for all }j\notin{\mathcal{I}}\right\},$
@@ 2.2 Prediction Model Architecture
where $\tilde{d}=\frac{d(d+1)}{2}+1$ and $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\tilde{d}}$ is defined as
$\phi(\bm{\theta})=\left[1,\bm{\theta}[1],\dots,\bm{\theta}[d],\frac{\bm{\theta}[1]^{2}-1}{\sqrt{2}},\dots,\frac{\bm{\theta}[d]^{2}-1}{\sqrt{2}},\bm{\theta}[1]\bm{\theta}[2],\dots,\bm{\theta}[d-1]\bm{\theta}[d]\right].$
@@
with parameters ${\bm{W}}_{B},{\bm{W}}_{C}\in\mathbb{R}^{d_{h}\times(\tilde{d}+1)},{\bm{w}}\in\mathbb{R}^{\tilde{d}},b\in\mathbb{R}$. Then, the $l$-th output can be expressed as
@@
To ensure a tractable optimization guarantee, we further introduce the following simplifications to our model:
@@ 2.3 Pretraining Algorithm
for $l=1,2$ and $T_{0}=0$. We employ a two-stage training procedure, as described in Algorithm 1, using these objectives.
@@
4 Initialize ${\bm{\gamma}}=(\gamma^{2},1,\dots,1,\gamma,\cdots,\gamma),{\bm{u}}(0)=m^{-1}{\bm{1}}_{m},{\bm{v}}(0)={\bm{1}}_{m},{\bm{a}}(0)={\bm{0}}_{m}$.
5 ${\bm{\gamma}}^{*}\leftarrow{\bm{\gamma}}^{(0)}-\eta\nabla_{{\bm{\gamma}}}\left(L_{1}({\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0))+\frac{\lambda_{1}}{2}\left\|{\bm{\gamma}}\right\|^{2}\right)$.
@@ 3.1 Preliminaries
  • The information exponent (Arous et al., 2021; Damian et al., 2023) of $h$ is defined as
@@
    $\mathrm{ge}(h):=\min_{{\mathcal{T}}\in L^{2}(P_{h})}\min\{i\in\mathbb{N}:H({\mathcal{T}}\circ h,i)\neq 0\},$
    where $L^{2}(P_{h})$ is the set of $L^{2}$-integrable functions with respect to $P_{h}$. Here, $P_{h}$ is the push-forward measure of the Gaussian measure by $h$.
@@
Previous works have established the sample complexity of various methods for learning a Gaussian single-index model. For example, kernel methods, which lack an adaptive basis, require at least $d^{\mathrm{deg}(g_{*})}$ samples (Ghorbani et al., 2021; Donhauser et al., 2021). In contrast, adaptive methods such as gradient-based methods on two-layer neural networks can achieve a sample complexity of ${\tilde{\mathcal{O}}}\left(d^{\Theta(\mathrm{ie}(g_{*}))}\right)$ by learning an adaptive feature map (Arous et al., 2021; Ba et al., 2022; Damian et al., 2022, 2023; Dandi et al., 2024). These approaches fall under the category of CSQ algorithms, and in this category, a sample complexity that depends on the information exponent is inevitable (Damian et al., 2022). However, recent works show that a nonlinear transformation introduced by data reuse (Arnaboldi et al., 2024; Lee et al., 2024) enables the algorithm to move into the broader class of Statistical Query (SQ) algorithms. This transformation allows the “effective” information exponent to be lowered to the generative exponent, thereby achieving a sample complexity of ${\tilde{\mathcal{O}}}\left(d^{\Theta(\mathrm{ge}(g_{*}))}\right)$.
@@ 3.2 Main Result
  • The context length is $N_{\mathrm{pt}}=\tilde{\Omega}\left(\max\left\{r^{3\mathrm{ge}(g_{*})}d^{8},T_{1}^{2}d^{4}\right\}\right)$.
@@
The conditions for the pretraining in our theorem can be satisfied with a pretraining sample complexity of $N_{\mathrm{pt}}=\Theta\left(d^{\Theta(\mathrm{ge}(g_{*}))}\right)$. In contrast, the pretraining sample complexities in previous works (Oko et al., 2024; Nishikawa et al., 2025) are governed by the information exponent, which can lead to a suboptimal rate in the worst case. This improvement is due to the nonlinearity of the MLP, as we discuss in detail in Section 4.
@@
Transformer Mamba
@@ 4 Proof Overview
In this section, we provide an overview of the proof for our theorem. The proof consists of three main parts: an analysis of one-step gradient descent on the Mamba layer, the optimization of the MLP, and a test error analysis. The formal proofs for each of these steps are provided in Appendices B, C, and D, respectively. In the following, we introduce the key ideas behind each step.
@@ 4.1 One-Step Gradient Descent on the Mamba Layer
${\bm{\gamma}}^{*}$ $\approx 2\eta\mathbb{E}_{({\bm{Z}},y)\sim{\mathcal{D}}(N_{\mathrm{pt}})}[y\nabla_{\bm{\gamma}}f({\bm{Z}};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0))]$
$\approx 2\eta\mathbb{E}_{\begin{subarray}{c}{\bm{\beta}}\sim\mathrm{Unif}(S_{r})\\ ({\bm{x}},y)\sim{\mathcal{D}}_{{\bm{\beta}}}\end{subarray}}\left[y\mathbbm{1}\left[\left\langle{\bm{c}}_{\bm{\beta}},{\bm{\gamma}}(0)\odot\phi({\bm{x}})\right\rangle>0\right]{\bm{c}}_{\bm{\beta}}\odot\phi({\bm{x}})\right],$
@@
In the absence of a gating mechanism and with only a linear attention, the term ${\bm{c}}_{\bm{\beta}}$ is replaced by $\mathbb{E}_{({\bm{x}},y)\sim{\mathcal{D}}_{\bm{\beta}}}[y\phi({\bm{x}})]$ and this term vanishes when $\mathrm{ie}(g_{*})>2$. This is a consequence of our input embedding using Hermite polynomials only up to the second degree, in combination with Stein’s lemma. As a result, pretraining in this case is unable to learn useful information. However, we prove that the nonlinear transformation introduced by the gating mechanism can reduce the information exponent to the generative exponent: $\mathrm{ie}(g_{*}\sigma(g_{*}/\rho+b))=\mathrm{ge}(g_{*})$. Combining Lemma 3.2 with our input embedding, we show that the gating enables the model to extract information. When $g_{*}$ is a non-even function, our result can also be shown to hold with the standard input embedding $\phi({\bm{x}})={\bm{x}}$, as can be seen from this intuition. Reducing the information exponent to the generative exponent crucially affects the achievement of a test-time sample complexity below the CSQ lower bounds. Nishikawa et al. (2025) shows that the softmax operator in the Transformer can also perform a nonlinear transformation on the label, which reduces the information exponent. This highlights a key difference in the mechanisms used by Mamba and the softmax Transformer to achieve this result.
@@
If the indicator $\mathbbm{1}[\cdot]$ inside the expectation is 1 with high probability, then the updated parameter $\bm{\gamma}$$^{*}$ can be approximated as ${\bm{\gamma}}^{*}\approx 2\eta\mathbb{E}_{{\bm{\beta}}\sim\mathrm{Unif}(S_{r})}\left[\mathbb{E}_{({\bm{x}},y\sim{\mathcal{D}}_{\bm{\beta}})}[y{\bm{c}}_{\bm{\beta}}\odot\phi({\bm{x}})]\right]$ and this close to zero when $\mathrm{ie}(g_{*})>2$. However, we show that the indicator deviates significantly from a constant value. We formally prove that this deviation allows the indicator to reduce the information exponent when multiplied by the label $y$, thereby inducing a sample complexity not governed by $\mathrm{ie}(g_{*})$. While Nishikawa et al. (2025) employ an architecture with a similar structure to ours—an MLP following a Softmax Transformer—they do not achieve the same improvement. This is because their use of Softmax places the model in a regime where a key indicator function is 1 with high probability. In addition, our observation for this improvement cannot be directly applied to the work of Oko et al. (2024) due to a key architectural difference: applying the MLP layer in the input embedding rather than at the output layer.
Pretrained Mamba Perform Test-Time Feature Learning.
We show that the pretrained parameter $\bm{\gamma}$$^{*}$ recovers the intrinsic feature space $S_{r}$ by attaining significantly larger components within the feature index set $\mathcal{I}$} than in other indices. Furthermore, we show that pretrained mamba performs test-time feature learning by establishing the following proposition:
Proposition 4.1 (Informal).
${\mathsf{Mamba}}({\bm{Z}};{\bm{\gamma}}^{*})\approx P_{1}+P_{2}\left(\frac{\langle{\bm{\beta}},{\bm{x}}\rangle}{r}\right)^{\mathrm{ge}(g_{*})}.$ (2)
@@ 4.2 Optimization of the MLP and Test Error Analysis
In our analysis of Stage II pretraining, we first construct an outer layer parameter $\bm{u}$$^{\prime}$ such that the loss $L_{2}({\bm{\gamma}}^{*},{\bm{u}}^{\prime},{\bm{v}},{\bm{a}})$ is sufficiently small and the norm $\left\|{\bm{u}}^{\prime}\right\|$ is well-bounded. Our construction is based on the techniques in Damian et al. (2022), which constructed a ReLU network approximating monomial transformations. This transformation allows our model to learn high-order polynomials with a few layers, in contrast to the multi-layer approach in Sun et al. (2025). More specifically, since Lemma 3.2 implies that $g_{*}(z)$ is a polynomial of $z^{\mathrm{ge}(g_{*})}$, and then we can construct an MLP layer approximating $g_{*}(\langle{\bm{\beta}},{\bm{x}}\rangle)$, when the input is provided in the form of (2).
@@ 5 Experiments
To support our theoretical findings, we pretrain and evaluate both Transformer and Mamba models on our data distribution. Our data is generated using a link function $g_{*}(z)=\mathrm{He}_{3}(z)/\sqrt{6}$ and an intrinsic dimension of $r=8$, with the varying ambient dimension $d=16,32$. We employ a 6-layer GPT-2 model (Radford et al., 2019) with 8 attention heads and a 12-layer Mamba model. To ensure a fair comparison, both models have an embedding dimension of 256 and a similar number of parameters. The overall experimental settings for pretraining follow those of prior works (Garg et al., 2022; Park et al., 2024). We also conduct kernel ridge regression on the intrinsic feature space to serve as a baseline for understanding the effect of feature learning. For this, we use a Gaussian RBF kernel with a bandwidth of 1 and a ridge parameter of 1. For evaluation, we measure the prediction error using squared error, with the number of context examples ranging from 1 to 40. We estimate the test error on 1024 randomly sampled tasks, using 2048 independent prompts for each task, and represent the results with the mean and standard deviation over these tasks. Figure 1 demonstrates our results. Both Transformer and Mamba models exhibit comparable performance that is rarely affected by the ambient dimension $d$. This suggests that both models mainly utilize information from the intrinsic feature space. In addition, these methods outperform kernel methods, even when we restrict the input of the kernel method to the intrinsic feature space. This observation aligns with our finding that Mamba, similar to Transformers, not only benefits from its adaptation to the intrinsic feature space but also performs test-time feature learning. [Uncaptioned image] Figure 1: Prediction error for in-context learning with Transformer and Mamba models, and kernel regression.

### v3

@@ 1.2 Related Works
Low-dimensional target function classes, such as sparse parities (Barak et al., 2022; Suzuki et al., 2023; Glasgow, 2024), signal-noise models (Allen-Zhu and Li, 2020; Cao et al., 2022), are widely adopted as theoretical benchmarks for studying a neural network’s ability to perform feature learning. This work specifically focuses on the single-index model. A line of theoretical work has analyzed the learning of these models and has established key results on sample complexity. The required sample complexity is governed by either the information exponent (for algorithms utilizing correlational information (Ben Arous et al., 2021; Bietti et al., 2022; Damian et al., 2023; Mousavi-Hosseini et al., 2023)) or the generative exponent (for algorithms that employ suitable label transformations (Damian et al., 2024; Lee et al., 2024; Arnaboldi et al., 2024; Joshi et al., 2024)). We discuss these sample complexity results in more detail in Section 3.1.
@@ 2.1 Data Distribution for In-Context Learning
    We draw the feature vector ${\bm{\beta}}\in\mathbb{R}^{d}$ uniformly from the unit sphere $S_{r}$ of a low-dimensional intrinsic feature space with dimension $r$, defined as:
    $S_{r}:=\left\{\bm{\theta}\in\mathbb{R}^{d}:\left\|\bm{\theta}\right\|=1,\bm{\theta}[j]=0\text{ for all }j\notin{\mathcal{I}}\right\},$
@@ 2.2 Prediction Model Architecture
where $\tilde{d}=1+d+\frac{d(d+1)}{2}$ and $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{\tilde{d}}$ is defined as
$\phi(\bm{\theta})=\left[1,\bm{\theta},\frac{\bm{\theta}\odot\bm{\theta}-{\bm{1}}_{d}}{\sqrt{2}},\bm{\theta}[1]\bm{\theta}[2],\dots,\bm{\theta}[d-1]\bm{\theta}[d]\right].$
@@
with parameters ${\bm{W}}_{B},{\bm{W}}_{C}\in\mathbb{R}^{d_{h}\times(\tilde{d}+1)},{\bm{w}}\in\mathbb{R}^{\tilde{d}+1},b\in\mathbb{R}$. Then, the $l$-th output can be expressed as
@@
To ensure a tractable optimization guarantee, we further introduce the following simplifications to our model, while preserving the essential mechanism:
@@ 2.3 Pretraining Algorithm
for $l=1,2$ and $T_{0}=0$. Using these objectives, we employ Algorithm 1 which proceeds in two stages: Stage I recovers intrinsic feature structures, enabling Stage II to perform in-context link estimation conditioned on these features. This approach enables tractable analysis and has been widely adopted in prior works (Ba et al., 2022; Damian et al., 2022).
@@
4 Initialize ${\bm{\gamma}}^{(0)}=(\gamma^{2},1,\dots,1,\gamma,\cdots,\gamma),{\bm{u}}^{(0)}=m^{-1}{\bm{1}}_{m},{\bm{v}}^{(0)}={\bm{1}}_{m},{\bm{a}}^{(0)}={\bm{0}}_{m}$.
5 ${\bm{\gamma}}^{*}\leftarrow{\bm{\gamma}}^{(0)}-\eta\nabla_{{\bm{\gamma}}}\left(L_{1}({\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)})+\frac{\lambda_{1}}{2}\left\|{\bm{\gamma}}\right\|^{2}\right)$.
@@ 3.1 Preliminaries
  • The information exponent (Ben Arous et al., 2021; Damian et al., 2023) of $h$ is defined as
@@
    $\mathrm{ge}(h):=\min_{{\mathcal{T}}\in L^{2}(\mathbb{P}_{h})}\min\{i\in\mathbb{N}:H({\mathcal{T}}\circ h,i)\neq 0\},$
    where $L^{2}(\mathbb{P}_{h})$ is the set of $L^{2}$-integrable functions with respect to $\mathbb{P}_{h}$. Here, $\mathbb{P}_{h}$ is the push-forward measure of the Gaussian measure by $h$.
@@
Previous works have established the sample complexity of various methods for learning a Gaussian single-index model. For example, kernel methods, which lack an adaptive basis, require at least $d^{\mathrm{deg}(g_{*})}$ samples (Ghorbani et al., 2021; Donhauser et al., 2021). In contrast, adaptive methods such as gradient-based methods on two-layer neural networks can achieve a sample complexity of ${\tilde{\mathcal{O}}}\left(d^{\Theta(\mathrm{ie}(g_{*}))}\right)$ by learning an adaptive feature map (Ben Arous et al., 2021; Ba et al., 2022; Damian et al., 2022, 2023; Dandi et al., 2024). These approaches fall under the category of CSQ algorithms, and in this category, a sample complexity that depends on the information exponent is inevitable (Damian et al., 2022). However, recent works show that a nonlinear transformation introduced by data reuse (Arnaboldi et al., 2024; Lee et al., 2024) enables the algorithm to move into the broader class of Statistical Query (SQ) algorithms. This transformation allows the “effective” information exponent to be lowered to the generative exponent, thereby achieving a sample complexity of ${\tilde{\mathcal{O}}}\left(d^{\Theta(\mathrm{ge}(g_{*}))}\right)$.
@@ 3.2 Main Result
  • The context length is $N_{\mathrm{pt}}=\tilde{\Omega}\left(r^{3\mathrm{ge}(g_{*})}d^{8}\vee T_{1}^{2}d^{4}\right)$.
@@
The conditions for the pretraining in our theorem can be satisfied with a pretraining sample complexity of $N_{\mathrm{pt}}=\Theta\left(d^{\Theta(\mathrm{ge}(g_{*}))}\right)$. In contrast, the pretraining sample complexities in previous works (Oko et al., 2024; Nishikawa et al., 2025) are governed by the information exponent, which can lead to a suboptimal rate in the worst case. This improvement is due to the nonlinearity of the MLP, as we discuss in detail in Section 4. Note that since these sample complexity bounds are sufficient conditions, it remains open whether Mamba is theoretically more efficient than Transformers in pretraining. While we leave this theoretical investigation as a future direction, our experiments in Section LABEL:sec:exp suggest this is indeed the case.
@@
Linear Transformer Softmax Transformer Mamba
@@ 4 Proof Overview
In this section, we provide an overview of the proof for our theorem. The proof consists of three main parts: an analysis of one-step gradient descent on the Mamba layer, the optimization of the MLP, and a test error analysis. The formal proofs for each of these steps are provided in Appendices B, C, and D, respectively. In the following, we introduce the key ideas behind each step, providing a detailed discussion of our theoretical components.
@@ 4.1 One-Step Gradient Descent on the Mamba Layer
We show that the pretrained parameter $\bm{\gamma}$$^{*}$ recovers the intrinsic feature space $S_{r}$ by attaining significantly larger components within the feature index set $\mathcal{I}$} than in other indices. Furthermore, we show that pretrained Mamba performs test-time feature learning by establishing the following proposition (formally stated in Proposition B.5):
Proposition 4.1 (Informal).
${\mathsf{Mamba}}({\bm{Z}};{\bm{\gamma}}^{*})\approx P_{1}+P_{2}\left(\frac{\langle{\bm{\beta}},{\bm{x}}\rangle}{r}\right)^{\mathrm{ge}(g_{*})},$ (2)
where $P_{1}$ and $P_{2}$ are independent of the data.
@@
${\bm{\gamma}}^{*}$ $\approx 2\eta\mathbb{E}_{({\bm{Z}},y)\sim{\mathcal{D}}(N_{\mathrm{pt}})}[y\nabla_{\bm{\gamma}}f({\bm{Z}};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)})]$
$\approx 2\eta\mathbb{E}_{\begin{subarray}{c}{\bm{\beta}}\sim\mathrm{Unif}(S_{r})\\ ({\bm{x}},y)\sim{\mathcal{D}}_{{\bm{\beta}}}\end{subarray}}\left[y\mathbbm{1}[\langle{\bm{c}}_{\bm{\beta}},{\bm{\gamma}}^{(0)}\odot\phi({\bm{x}})\right\rangle>0]{\bm{c}}_{\bm{\beta}}\odot\phi({\bm{x}})],$
@@
For the proof of Proposition 4.1 and our test-time sample complexity, nonlinear transformation introduced by gating mechanism plays a crucial role. In the absence of a gating mechanism and with only a linear attention, the term ${\bm{c}}_{\bm{\beta}}$ is replaced by $\mathbb{E}_{({\bm{x}},y)\sim{\mathcal{D}}_{\bm{\beta}}}[y\phi({\bm{x}})]$ and this term vanishes when $\mathrm{ie}(g_{*})>2$. This is a consequence of our input embedding using Hermite polynomials only up to the second degree, in combination with Stein’s lemma. As a result, pretraining in this case is unable to learn useful information. However, we prove that the gating mechanism ensures $c_{\bm{\beta}}$ is non-zero, thereby enabling the model to extract information. This is because the nonlinear transformation introduced by the gating mechanism reduces the information exponent to the generative exponent: $\mathrm{ie}(g_{*}\sigma(g_{*}/\rho+b))=\mathrm{ge}(g_{*})$. This reduction, combined with Lemma 3.2 and our input embedding, makes information extraction possible. When $g_{*}$ is a non-even function, our result can also be shown to hold with the standard input embedding $\phi({\bm{x}})={\bm{x}}$, as can be seen from this intuition. Reducing the information exponent to the generative exponent crucially affects the achievement of a test-time sample complexity below the CSQ lower bounds. Nishikawa et al. (2025) shows that the softmax operator in the Transformer can also perform a nonlinear transformation on the label, which reduces the information exponent. This highlights a key difference in the mechanisms used by Mamba and the softmax Transformer to achieve this result.
@@
The nonlinearity of the MLP is crucial for our analysis of the pretraining sample complexity. If the indicator $\mathbbm{1}[\cdot]$ inside the expectation is 1 with high probability, then the updated parameter $\bm{\gamma}$$^{*}$ can be approximated as ${\bm{\gamma}}^{*}\approx 2\eta\mathbb{E}_{{\bm{\beta}}\sim\mathrm{Unif}(S_{r})}\left[\mathbb{E}_{({\bm{x}},y\sim{\mathcal{D}}_{\bm{\beta}})}[y{\bm{c}}_{\bm{\beta}}\odot\phi({\bm{x}})]\right]$ and this close to zero when $\mathrm{ie}(g_{*})>2$, leading to less information gain. However, we show that the indicator deviates significantly from a constant value. We formally prove that this deviation allows the indicator to reduce the information exponent when multiplied by the label $y$, thereby inducing a pretraining sample complexity not governed by $\mathrm{ie}(g_{*})$. While Nishikawa et al. (2025) employ an architecture with a similar structure to ours—an MLP following a Softmax Transformer—they do not achieve the same improvement. This is because their use of Softmax places the model in a regime where a key indicator function is 1 with high probability, which is sufficient in their case as Softmax generates the necessary higher-order functions of the input, while our analysis only uses up to second-order functions. In addition, our observation for this improvement cannot be directly applied to the work of Oko et al. (2024) due to a key architectural difference: applying the MLP layer in the input embedding rather than at the output layer.
@@ 4.2 Optimization of the MLP and Test Error Analysis
In our analysis of Stage II pretraining, we show that the MLP can fit the link function $g_{*}$. We first construct an outer layer parameter $\bm{u}$$^{\prime}$ such that the loss $L_{2}({\bm{\gamma}}^{*},{\bm{u}}^{\prime},{\bm{v}},{\bm{a}})$ is sufficiently small and the norm $\left\|{\bm{u}}^{\prime}\right\|$ is well-bounded. Our construction is based on the techniques in Damian et al. (2022), which constructed a ReLU network approximating monomials. This allows our model to learn high-order polynomials with a few layers, in contrast to the multi-layer approach in Sun et al. (2025). This is possible because Lemma 3.2 implies that $g_{*}(z)$ is a polynomial of $z^{\mathrm{ge}(g_{*})}$, allowing us to construct an MLP that approximates $g_{*}(\langle{\bm{\beta}},{\bm{x}}\rangle)$, when the input is provided in the form of (2).
@@ 5 Experiments
In this section, we present experimental results on standard architectures and optimization algorithms beyond our simplified setting to support our theoretical findings.
We pretrain and evaluate both Transformer and Mamba models on our data distribution. Our base configuration uses a link function $g_{*}(z)=\mathrm{He}_{3}(z)/\sqrt{6}$, an intrinsic dimension of $r=8$, and an ambient dimension $d=32$. We employ a 6-layer GPT-2 model (Radford et al., 2019) with 8 attention heads and a 12-layer Mamba model. To ensure a fair comparison, both models have an embedding dimension of 256 and a similar number of parameters. The overall experimental settings for pretraining follow those of prior works (Garg et al., 2022; Park et al., 2024), including the optimization algorithm and a pretraining context length $N_{\mathrm{pt}}=64$. We also conduct kernel ridge regression on the intrinsic feature space to serve as a baseline for understanding the effect of feature learning. For this, we use a Gaussian RBF kernel with a bandwidth of 1 and a ridge parameter of 1. For evaluation, we measure the prediction error using squared error, with the number of context examples ranging from 1 to 40. We estimate the test error on 1024 randomly sampled tasks, using 2048 independent prompts for each task, and represent the results with the mean and standard deviation over these tasks.
To validate our theoretical results on how problem parameters influence performance, we then analyze trends by varying parameters from our base configuration: the ambient dimension to $d=16$, the intrinsic dimension to $r=16$, and the pretraining context length to $N_{\mathrm{pt}}=16$. Figure 1(a) demonstrates the influence of the ambient dimension $d$. Both Transformer and Mamba models exhibit comparable performance that is rarely affected by the ambient dimension $d$. In contrast, when the intrinsic dimension $r$ is increased, both models exhibit performance degradation, while their performance remains comparable (Figure 1(b)). This suggests that both models mainly utilize information from the intrinsic feature space. In addition, these methods outperform kernel methods, even when we restrict the input of the kernel method to the intrinsic feature space. This observation aligns with our finding that Mamba, similar to Transformers, not only benefits from its adaptation to the intrinsic feature space but also performs test-time feature learning. Lastly, when using a small pretraining context length $N_{\mathrm{pt}}=16$, the Transformer’s performance deteriorates significantly, while Mamba’s performance remains comparable to its performance with $N_{\mathrm{pt}}=64$ (Figure 1(c)). This observation aligns with our pretraining sample complexity result, which is lower than that of the Transformer established by Nishikawa et al. (2025).
(a) $d=32$ vs. $d=16$
(b) $r=8$ vs. $r=16$
(c) $N_{\mathrm{pt}}=64$ vs. $N_{\mathrm{pt}}=16$
Figure 1: Comparison of prediction error for in-context learning with Transformer and Mamba models, and kernel regression across different problem parameter settings.